OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of ? y = -5x + 22 y = 5x + 22

OpenStudy (owlcoffee):

You missed the slope there. But let's take a closer look to the slope formula: \[m=\frac{ y_2 - y_1 }{ x_2 - x_1 }\] So we know that the way to calculate the slope is the division between the change in y and the change in x. Buw if we know the slope and one of the points that belong to that line then I must be able to find the equation of the line by making it have that structure, by that, let's multiply both sides by (x2 - x1) : \[m( x_2 - x_1 ) = ( y_2 - y_1 )\] So, this must mean that at least one of them must be generic, in order to find the line that belong to any point (x,y) and a known point (a,b) I'd just have to replace: \[m(x-a)=(y-b)\] Doing the distributive: \[mx-ma=y-b\] \[mx-y+b-ma=0\] We can see it has the form "Ax+By+C=0" wich corresponds to a line, giving me the certainty that this is a valid formula to find the line knowing a point an it's slope. We conclude that the formula is: \[(y- y_o)=m(x- x_o )\] Where "m" is the slope and "xo" and "yo" are the coordinates of a known point.